# Deep Neural Networks: Multi-Classification and Universal Approximation

### v1

@@ 1.1. Problem formulation
Let us consider $1\leq p<\infty$, two integers $m,\,d\geq 1$, and a bounded domain $\Omega\subset\mathbb{R}^{d}$. Then, for any $f\in L^{p}(\Omega;\mathbb{R}^{m}_{+})$ and $\varepsilon>0$, there exist $\mathcal{L}=\mathcal{L}(\varepsilon)\geq 1$ and parameters $\mathcal{W}^{\mathcal{L}},$ and $\mathcal{B}^{\mathcal{L}}$ such that the input-output map of (1.1) with a width of $w_{max}=\max\{d+1,2m\}$, satisfies
@@ 1.3. Related Work
Deep learning has gained popularity due to its state-of-the-art performance in various machine learning applications [26, 40]. In practice, neural networks are typically trained using optimization methods minimizing a least-squares error functional, with stochastic gradient descent algorithms serving as an essential tool to search for minimizers. While this numerical approach, combined with backpropagation techniques to compute the gradients, often leads to solutions that outperform human experts, we still lack a solid mathematical understanding of why deep learning works so well. The results in this paper aim to contribute to explain such performance by explicit constructions, which yield concrete estimates of the complexity required for neural networks to achieve the desired goals, memorization, and universal approximation.
In this section, we delve into recent advancements in this context.
@@
In the context of deep neural networks, one of the first attempts for sigmoid functions was presented in [47]. For neural networks with ReLU activation functions, [48] demonstrated that for a fixed depth $L$ and $N$ points, a neural network with a width depending on $N$ and satisfying a technical assumption can memorize the data. In particular, this result holds if there exists $l>1$ such that $d_{j}d_{j+l}=O(Nm)$ for some $j>1$. In [34], it is shown that ReLU networks with a width greater than $3$ and $O(N^{2/3}log(M))$ neurons suffice to approximately memorize $N$ points with $M$ one-dimensional classes, in the sense that the data can be driven to be $\varepsilon$-close to the labels, where $\varepsilon>0$ (this may be referred to as approximate simultaneous controllability). The same article also establishes that a 3-wide neural network with $O(N^{2/3}log(M))$ layers (or neurons) is sufficient for such approximate memorization. Finally, [45] showed that, by fixing a width of $12$ neurons, networks can memorize any $N$ points with $M$ one-dimensional classes using $O(N^{1/2}+log(M))$ layers (or neurons). See [45] for the exact expression of the required depth.
From the control theory perspective, in [37, 3], simultaneous controllability for ResNets and neural ordinary differential equations is proven (see also [1]). This implies the memorization property. In particular, it is shown that it is possible to memorize a $d-$dimensional dataset of $N$ points and $M$ classes using a ReLU ResNet with fixed width $d$ and $O(N)$ layers and, therefore, $O(dN)$ neurons. The novelty in [37] lies in the genuinely constructive approach to building parameters. There exists an extensive literature encompassing the controllability of linear and nonlinear dynamics in which constructive methods are employed, independent of the learning objectives, which constitute our motivation (see [6, 46, 15, 24, 25] and the references therein). An extension of [37] can be found in [8], where it is shown that for neural networks with sufficiently large depth but fixed width, interpolation can be guaranteed through the use of non-linear activation functions.
Universal Approximation Theorem: Classical results in this field primarily focus on shallow neural networks [4, 10, 20, 30, 36], with Cybenko’s celebrated work [10] as a notable example, who proved that a single hidden layer neural network could approximate any continuous function within a compact set of $\mathbb{R}^{n}$ using a sigmoidal activation function. However, such density results trace back to 1932, with Wiener’s Tauberian theorem, which provides necessary and sufficient conditions under which any function in $L^{1}(\mathbb{R})$ or $L^{2}(\mathbb{R})$ can be approximated by linear combinations of translations of a given $L^{1}(\mathbb{R})$ function.
@@
With regard to universal approximation in $L^{p}$ spaces, [33] demonstrates that a deep neural network with ReLU activation function and a width of $d+4$ can approximate any function in $L^{1}(\mathbb{R}^{d};\mathbb{R})$. They allocate $d$ neurons for transferring input information to subsequent layers, two neurons to carry the information of the approximation made by the previous layers, and two neurons for approximation on each layer. The same article also proves that if the width of a deep neural network is less than or equal to $d$, it is impossible to approximate $L^{1}(\mathbb{R}^{d};\mathbb{R})$ or $L^{1}(\Omega;\mathbb{R})$ for a compact $\Omega$. In [27], for $p\in[1,\infty)$, a compact set $\Omega\subset\mathbb{R}^{d}$, and $m\geq 1$, it is established that it is possible to approximate the spaces $L^{p}(\mathbb{R}^{d};\mathbb{R}^{m})$ and $L^{p}(\Omega;\mathbb{R}^{m})$ with a ReLU network of width $d+m+1$. Their main argument for approximating $L^{p}(\mathbb{R}^{d};\mathbb{R}^{m})$ involves using a neural network to approximate cutoff functions. For $L^{p}(\Omega;\mathbb{R}^{m})$ functions, they prove universal approximation for $C(\Omega;\mathbb{R}^{m})$ functions using $m+d+1$ neurons, concluding by density. More precise estimates of the minimal width in $L^{p}(\mathbb{R}^{d};\mathbb{R}^{m})$ and $L^{p}(\Omega;\mathbb{R}^{m})$ are presented in [35], which determine the minimal widths to be $\max\{d+1,m\}$ and $\max\{d+2,m+1\}$, respectively. The proof of this theorem utilizes a coding scheme, consisting of encoding (projecting) $x\in\Omega$ into a codeword (scalar values) containing information about $x$, and then a decoder transforming each codeword into a target function $f(x)$. This scheme is applied to approximate continuous functions and completed with density arguments in $L^{p}$. In the particular case of the Leaky-ReLU activation function, a variant of ReLU, it was proven in [7] that for a compact $\Omega\subset\mathbb{R}^{d}$, the minimum width required to approximate functions in $L^{p}(\Omega;\mathbb{R}^{m})$ is $\max\{d,m,2\}$. Recently, [28] has shown that the minimum width of a neural network with a ReLU activation function necessary to approximate $L^{p}(\Omega;\mathbb{R}^{m})$ is $\max\{d,m,2\}$, with $\Omega$ being a compact set. The proof of this result, based on [35], employs the coding scheme to approximate continuous functions in a compact $\Omega$, concluding the result for $L^{p}(\Omega;\mathbb{R}^{m})$ functions by density.
Universal approximation theorems for the space of continuous functions in the case of arbitrary depth are discussed in [16, 17, 18, 27, 35, 42, 31]. In particular, the minimal width for approximating $L^{p}$ functions using the ResNet architecture has been studied in [41]. For results of universal approximation theorems using transforms to approximate $L^{p}$ functions, see [2]. For an extended introduction to the universal approximation theorem in more general spaces, see the survey article [13].
@@
In Theorem 1.1, we prove that the neural network defined by (1.1) satisfies the finite sample memorization property with a width no greater than $2$ and depth $L=2N+4M-1$. Consequently, we use $O(N)$ neurons. To the best of our knowledge, this is the first result for $2-$wide deep neural networks. Furthermore, our geometric analysis, utilized to prove finite sample memorization, allows us to develop a new and constructive method to explore the behavior of networks as in (1.1). While our construction is motivated by [37], the strategy followed is different because the dynamic behavior of (1.1) differs significantly from the ResNet. Indeed, while the activation function in ResNet merely acts as an infinitesimal increment, here we deal with a genuine time-discrete non-linear dynamics.
Regarding universal approximation, our primary contribution lies in our purely constructive proof. Unlike other constructive proofs [35, 13, 28], ours provides a detailed geometric step-by-step description. This approach allows for visualization and geometric interpretation of the neural network’s operations at each layer, akin to [37], demonstrating how the network approximates simple functions supported on hyperrectangles. Consequently, it enables approximation of entire function spaces $L^{p}(\Omega;\mathbb{R}_{+})$ and $L^{p}(\Omega;\mathbb{R})$. Furthermore, this explicit construction enables us to provide the first estimates of the required depth for a neural network of width $d+1$ to approximate the function in terms of its $W^{1,p}$ norm.
@@ 1.5. Outline
The rest of the paper is organized as follows: In Section 2, we conduct a geometric analysis of the discrete system (1.1), introducing fundamental tools essential to our proofs. In Section 3, we offer an informal demonstration on constructing parameters to guarantee the finite sample memorization property, illustrated with a specific example. Section 4 contains the formal proof of Theorem 1.1, followed by the proof of the universal approximation theorem (Theorem 1.2). Finally, in Section 6, we discuss extensions and open problems.
@@ 2.2. Projection Lemma
We begin by observing that (2.3) is equivalent to $v\cdot(x_{i}-x_{j})\neq 0$, for every $i\neq j\in\llbracket 1,N\rrbracket$. For each $i,j$, we have $v\cdot(x_{i}-x_{j})=\|x_{i}-x_{j}\|_{\mathbb{R}^{d}}\cos(\theta(i,j)),$ where $\theta(i,j)\in[0,2\pi)$ is the angle between the vector $v$ and $x_{i}-x_{j}$. Therefore,
$v\cdot(x_{i}-x_{j})\neq 0,\quad\Longleftrightarrow\quad\theta(i,j)\notin\left\{\frac{\pi}{2},\frac{3\pi}{2}\right\}.$ (2.4)
Let us consider the sets
$Q:=\left\{q(i,j)=x_{i}-x_{j}\,:\,i\neq j\in\llbracket 1,N\rrbracket\right\},$
and
$G:=\left\{v\in\mathbb{S}^{d-1}\,:\,v\cdot q(i,j)=0,\text{ for some }i\neq j\in\llbracket 1,N\rrbracket\right\}.$
We must check that $\mathbb{S}^{d-1}\setminus G\neq\emptyset.$ Note that the cardinality of $Q$ corresponds to the number of interconnections between the points $\{x_{i}\}_{i=1}^{N}$, and it is equal to $(N-1)N/2$. From (2.4), for each $q(i,j)\in Q$, there exist just two angles such that $v\cdot q(i,j)=0$. Consequently, we have that $|G|=(N-1)N$ and therefore $G^{c}\neq\emptyset$. ∎
@@
  1. (1)
    The results remain true for a countable set of data to be projected, but it could fail for a continuous set.
  2. (2)
    When the dataset is finite, $G^{c}$ is non-empty and contains all vectors of $\mathbb{S}^{d-1}$ except for a finite number of them. Hence, for each $g\in G$ and every $\varepsilon$ there exists $g_{\varepsilon}$ in an $\varepsilon-$neighborhood of $g$ such that $g_{\varepsilon}\in G^{c}$. In other words, the stated projection property is open.
  3. (3)
    This result also holds true when we consider any finite dataset, and $v\in S\subset\mathbb{S}^{d-1}$, where $S$ is an infinite countable set.
@@
At this point, it becomes evident that at least two hyperplanes are necessary to develop an algorithm for classifying data. Indeed, if we were restricted to using just a single hyperplane, we would be unable to develop a compression process for any data set. Therefore, it is not feasible to classify any $d-$dimensional dataset using a $1$-wide neural network. Thus, $2$ is the minimal width for a neural network (as defined by (1.1)) capable of addressing any classification problem.
@@ 4.1. Preconditioning of the data:
$w_{1}\cdot x_{i}+b_{1}>0,\quad\forall i\in\llbracket 1,N\rrbracket$
@@ 4.2.1. The induction strategy
In the sequel we write $\phi(\mathcal{W}^{L},\mathcal{B}^{L},\mathcal{C}_{k})=z_{k}$ when $\phi(\mathcal{W}^{L},\mathcal{B}^{L},x)=z$ for every $x\in\mathcal{C}_{k}$, whenever all points corresponding to a class $\mathcal{C}_{k}$ collapse on the same point. Therefore, compressing all the classes is equivalent to proving the existence of parameters $\mathcal{W}^{\hat{L}}$ and $\mathcal{B}^{\hat{L}}$, $\hat{L}>0$, and a sequence of different vectors $\{z_{k}\}_{k=0}^{M}\subset\mathbb{R}^{2}$ such that
$\phi^{\hat{L}}(\mathcal{W}^{\hat{L}},\mathcal{B}^{\hat{L}},\mathcal{C}_{k})=z_{k},\qquad\text{for every }k\in\llbracket 0,M-1\rrbracket.$ (4.4)
@@
Let $\{x_{i},y_{i}\}_{i=1}^{N}\subset\mathbb{R}\times\llbracket 0,M-1\rrbracket$ and assume that $x_{i}\neq x_{j}$ for every $i\neq j\in\llbracket 1,N\rrbracket$. For every $k\in\llbracket 0,M-1\rrbracket$ we assume that there exist $z_{0}\in\mathbb{R}^{2}$, $\tilde{L}\geq 1$, $\mathcal{W}^{\tilde{L}}$ and $\mathcal{B}^{\tilde{L}}$ such that
$\phi(\mathcal{W}^{\tilde{L}},\mathcal{B}^{\tilde{L}},\mathcal{C}_{k})=z_{0},\qquad\phi(\mathcal{W}^{\tilde{L}},\mathcal{B}^{\tilde{L}},\mathcal{C}\setminus\mathcal{C}_{k})\neq z_{0},$ (4.5)
@@
$\phi(\mathcal{W}^{\tilde{L}},\mathcal{B}^{\tilde{L}},z^{1})\neq\phi^{\hat{L}}(\mathcal{W}^{\tilde{L}},\mathcal{B}^{\tilde{L}},z^{2}),\quad\text{for all }z^{1},\,z^{2}\in\mathcal{C}\setminus\mathcal{C}_{k},\,z^{1}\neq z^{2}.$ (4.6)
@@
$\phi^{L_{1}}(\mathcal{W}^{L_{1}},\mathcal{B}^{L_{1}},\mathcal{C}_{k})=z_{k},$
@@
In other words, Proposition 4.1 demonstrates that to compress all the classes of points, it is sufficient to compress a single (but arbitrary) class of points without collapsing the points not belonging to that class.
@@
  • Data structuring: Construction of hyperplanes driving the data set to some particular structure.
  • Compression process: Using the structure established in the prior step, introduce hyperplanes to collapse points belonging to the same class.
@@
$b_{3}^{1}=-\left(\frac{a^{2}_{r_{1}-1}+a^{2}_{r_{1}}}{2}\right),\quad b_{3}^{2}=\frac{a^{2}_{r_{1}}+a^{2}_{r_{1}+2}}{2}$ (4.9)
@@
(2) Inductive Step: The initial step has been achieved. We now aim to show that induction can also be applied successfully. In fact, to do that, it suffices to apply the arguments of the initial step again and again. Note that the model under consideration is, in this respect, simpler to handle than, for instance, ResNets, since in the present one, points collapse to the exact same location, and once this happens, they will never split again in the forthcoming iterations. In this way, if $\alpha_{k}$ denotes the number of elements in $\mathcal{C}_{k}$ i.e. $|\mathcal{C}_{k}|=\alpha_{k}$, all points in $\mathcal{C}_{k}$ will collapse applying $2\alpha_{k}$ times the nonlinear mapping $\sigma$, so that $L_{k}=2\alpha_{k}$.
@@
Data structuring: Denote by $x_{r_{t}}^{2j+1}$ the elements of $\mathcal{C}_{k}^{2j+1}$ for $t\in\llbracket 1,\alpha_{k}\rrbracket$. Observe that, after compressing the first $j$ elements of $\mathcal{C}_{k}^{2j+1}$, we will always have that $x_{r_{t}}^{2j+1}=(0,0)$ for $t\in\llbracket 1,j\rrbracket$ and $x_{r_{j+1}}^{2j+1}=(a_{j+1}^{2j+1},0)$ for some $a_{j+1}^{2j+1}\in\mathbb{R}$. Since we assumed from the beginning that they are not neighboring points of the same class, let $x_{s}^{2j+1}=(a^{2j+1}_{s},0)\notin\mathcal{C}_{k}^{2j+1}$ be the nonzero closest point to $x_{r_{1}}^{j}$ in the $x-$axis. We also consider $x_{s+1}^{2j+1}=(a_{s+1}^{2j+1},0)$ with $a_{s+1}^{2j+1}\neq 0$, the point to the right of $x_{s}^{2j+1}$ on the $x$-axis (if it does not exist, we take $x_{s+1}^{2j+1}:=x_{s}^{2j+1}+(0.5,0)$).
@@
Compression: We are going to compress $x_{r_{j}}^{2j+2}$ with $x_{r_{j+1}}^{2j+2}$. Let $x_{u}^{2j+2}=(0,a_{u}^{2j+2})$ and $x_{d}^{2j+2}=(0,a_{d}^{2j+2})$ be the points lying above and below $x_{r_{j}}^{2j+2}$ on the $y-$axis, respectively. Also, denote by $x_{rr}^{2j+2}=(a_{rr}^{2j+2},0)$ and $x_{ll}^{2j+2}=(a_{ll}^{2j+2},0)$ the points that are to the right and left of $x_{j+1}^{2j+2}$ in the $x-$axis. Thus, we consider the parameters

### v2

@@ 1.3.2. Implications for Neural Network Training
Let us consider $1\leq p<\infty$, two integers $m,\,d\geq 1$, and a bounded domain $\Omega\subset\mathbb{R}^{d}$. Then, for any $f\in L^{p}(\Omega;\mathbb{R}^{m}_{+})$ and $\varepsilon>0$, there exist $\mathcal{L}=\mathcal{L}(\varepsilon)\geq 1$ and parameters $\mathcal{W}^{\mathcal{L}},$ and $\mathcal{B}^{\mathcal{L}}$ such that the input-output map of (1.2) with width $w_{max}=\max\{d+1,2m\}$, satisfies
@@ 1.4. Related Work
Deep learning has gained popularity due to its state-of-the-art performance in various machine learning applications [24, 38]. In practice, neural networks are typically trained using optimization methods minimizing a least-squares error functional, with stochastic gradient descent algorithms serving as an essential tool to search for minimizers. While this numerical approach, combined with backpropagation techniques to compute the gradients, often leads to solutions that outperform human experts, we still lack a solid mathematical understanding of why deep learning works so well. The results in this paper aim to contribute to explain such performance by explicit constructions, which yield concrete estimates of the complexity required for neural networks to achieve the desired goals – in particular memorization, and universal approximation – and explicit bounds for the trained parameters.
In this section, we describe some recent related advancements in this broad field.
@@
In the context of deep neural networks, [45] constitutes one of the first attempts with sigmoid activation functions. For the ReLU activation function, [46] demonstrated that for a fixed depth $L$, a neural network with a width depending on $N$ and satisfying a technical assumption can memorize $N$ data. In particular, this result holds if there exists $l>1$ such that $d_{j}d_{j}+l=O(Nm)$ for some $j>1$. In [32], it is shown that ReLU networks with a width greater than $3$ and $O(N^{2/3}log(M))$ neurons suffice to approximately memorize $N$ points with $M$ one-dimensional classes, in the sense that the data can be driven to be $\varepsilon$-close to the labels, where $\varepsilon>0$ (this constitutes an approximate simultaneous or ensemble controllability result). The same article also establishes that a 3-wide neural network with $O(N^{2/3}log(M))$ layers (or neurons) suffices for such approximate memorization. Finally, [43] showed that, by fixing a width $12$, networks can memorize any $N$ points with $M$ one-dimensional classes using $O(N^{1/2}+log(M))$ layers (or neurons). See [43] for the exact expression of the required depth. We highlight that none of the previously mentioned works provides an explicit estimate of the parameter norms ($\ell^{2}$ or $\ell^{\infty}$); they focus mainly on the number of parameters.
From the control theory perspective, in [35, 3], simultaneous controllability for ResNets and neural ordinary differential equations is proven (see also [1]). This implies the memorization property. The novelty in [35] lies in the genuinely constructive approach to building parameters. An extension of [35] can be found in [8], where it is shown that for neural networks with sufficiently large depth but fixed width, interpolation can be guaranteed through the use of non-linear activation functions. Finally, we mention that constructive methods to prove controllability have been widely used in classical control theory; see [6, 44, 14, 22, 23] and the references therein for a detailed discussion.
Universal Approximation Theorem: Classical results in this field primarily focus on shallow neural networks [4, 10, 18, 28, 34], with Cybenko’s celebrated work [10] as a notable example, who proved that a single hidden layer neural network could approximate any continuous function within a compact set of $\mathbb{R}^{n}$ using a sigmoidal activation function. However, such density results trace back to 1932, with Wiener’s Tauberian theorem, which provides necessary and sufficient conditions under which any function in $L^{1}(\mathbb{R})$ or $L^{2}(\mathbb{R})$ can be approximated by linear combinations of translations of a given $L^{1}(\mathbb{R})$ profile, the gaussian, for instance.
@@
With regard to universal approximation in $L^{p}$ spaces, [31] demonstrates that a deep neural network with the ReLU activation function and width $d+4$ can approximate any function in $L^{1}(\mathbb{R}^{d};\mathbb{R})$. They allocate $d$ neurons for transferring input information to subsequent layers, two neurons to carry the information of the approximation made by the previous layers, and two neurons for approximation on each layer. The same article also proves that if the width of a deep neural network is less than $d$, it is impossible to approximate $L^{1}(\mathbb{R}^{d};\mathbb{R})$ or $L^{1}(\Omega;\mathbb{R})$ for a compact $\Omega$. In [25], for $p\in[1,\infty)$, a compact set $\Omega\subset\mathbb{R}^{d}$, and $m\geq 1$, it is established that it is possible to approximate the spaces $L^{p}(\mathbb{R}^{d};\mathbb{R}^{m})$ and $L^{p}(\Omega;\mathbb{R}^{m})$ with a ReLU network of width $d+m+1$. Their main argument for approximating $L^{p}(\mathbb{R}^{d};\mathbb{R}^{m})$ involves using a neural network to approximate cutoff functions. For $L^{p}(\Omega;\mathbb{R}^{m})$ functions, they prove universal approximation for $C(\Omega;\mathbb{R}^{m})$ functions using $m+d+1$ neurons, concluding by density. More precise estimates of the minimal width in $L^{p}(\mathbb{R}^{d};\mathbb{R}^{m})$ and $L^{p}(\Omega;\mathbb{R}^{m})$ are presented in [33], which determine the minimal widths to be $\max\{d+1,m\}$ and $\max\{d+2,m+1\}$, respectively. The proof of this theorem utilizes a coding scheme, consisting of encoding (projecting) $x\in\Omega$ into a codeword (scalar values) containing information about $x$, and then a decoder transforming each codeword into a target function $f(x)$. This scheme is applied to approximate continuous functions and completed with density arguments in $L^{p}$. In the particular case of the Leaky-ReLU activation function, a variant of ReLU, in [7] it was proven that for a compact $\Omega\subset\mathbb{R}^{d}$, the minimum width required to approximate functions in $L^{p}(\Omega;\mathbb{R}^{m})$ is $\max\{d,m,2\}$. Recently, [26] has shown that the minimum width of a neural network with a ReLU activation function necessary to approximate $L^{p}(\Omega;\mathbb{R}^{m})$ is $\max\{d,m,2\}$, $\Omega$ being a compact set. The proof of this result, based on [33], employs the coding scheme to approximate continuous functions in a compact $\Omega$, concluding the result for $L^{p}(\Omega;\mathbb{R}^{m})$ functions by density.
Universal approximation theorems for the space of continuous functions in the case of arbitrary depth are discussed in [15, 16, 25, 33, 40, 29]. In particular, the minimal width for approximating $L^{p}$ functions using the recurrent neural networks (RNN) architecture has been studied in [39]. We refer to [2] for universal approximation theorems in $L^{p}$, using transformers. For an extended introduction to the universal approximation theorem in more general spaces, see the survey article [12].
@@
In Theorem 1.1, we present a constructive proof of the fact that the neural network defined by (1.2) satisfies the finite sample memorization property with width $2$ and depth $L=2N+4M-1$, which implies a total number of neurons of order $O(N)$. To the best of our knowledge, this is the first purely constructive and geometrically interpretable proof of memorization for narrow MLPs. Our approach allows not only for an explicit step-by-step construction of the network parameters but also provides explicit upper bounds on their norms. This has important implications for regularized training: minimizing a standard empirical loss with $\ell^{2}$-regularization leads to parameters whose norms are uniformly bounded by the ones of our constructed parameters. In the limit of vanishing regularization, this guarantees exact interpolation while maintaining control over the norm of the trained parameters.
Regarding universal approximation, for MLP of fixed width $d+1$, we establish universal approximation results in $L^{p}(\Omega;\mathbb{R}+)$ and $L^{p}(\Omega;\mathbb{R})$, for any $p\in[1,\infty)$ and bounded domain $\Omega\subset\mathbb{R}^{d}$. Again, our primary contribution lies in our purely constructive proof based on a detailed geometric and recursive argument. In contrast to other constructive results such as [33, 12, 26], our approach emphasizes geometric intuition and visualization at each layer, following ideas similar to [35]. While the width requirement in [26] is optimal (namely, width equals $d$), our construction requires width $d+1$ but yields a fully interpretable and elementary strategy to understand how approximation is achieved. Moreover, all parameters involved are given explicitly, which allows us to estimate the depth required to approximate functions with a prescribed accuracy.
@@ 1.6. Outline
The rest of the paper is organized as follows: In Section 2, we conduct a geometric analysis of the discrete system (1.2), introducing fundamental tools essential to our proofs. In Section 3, we offer an informal demonstration on constructing parameters to guarantee the finite sample memorization property, illustrated with a specific example. Section 4 contains the formal proof of Theorem 1.1, followed by the proof of the universal approximation theorem (Theorem 1.4). Section 6 provides the proof of the theorems related to the regularized training and vanishing regularization. Finally, in Section 7, we discuss extensions and open problems.
@@ 2.2. Projection Lemma
For each pair $1\leq i<j\leq N$, define the set of directions
$H_{ij}=\big{\{}v\in\mathbb{S}^{d-1}\,:\,v\cdot(x_{i}-x_{j})=0\big{\}}.$
Since $x_{i}\neq x_{j}$, the vector $q_{ij}:=x_{i}-x_{j}\in\mathbb{R}^{d}\setminus\{0\}$, so $H_{ij}$ is the intersection of the unit sphere with the hyperplane orthogonal to $q_{ij}$. Hence, $H_{ij}$ is a closed subsphere of dimension $d-2$ and thus a proper, nowhere‐dense subset of $\mathbb{S}^{d-1}$. Set
$G=\bigcup_{1\leq i<j\leq N}H_{ij}.$
Then $G$ is a finite union of closed, nowhere‐dense subsets of the compact manifold $\mathbb{S}^{d-1}$. By the Baire Category Theorem (or by observing that each $H_{ij}$ has surface‐measure zero and hence their finite union cannot exhaust the full‐measure sphere), we conclude $\mathbb{S}^{d-1}\setminus G\neq\emptyset$. Any $v\in\mathbb{S}^{d-1}\setminus G$ satisfies
$v\cdot(x_{i}-x_{j})\neq 0,\quad\forall i\neq j,$
which is equivalent to $v\cdot x_{i}\neq v\cdot x_{j}$ whenever $i\neq j$, concluding the proof. ∎
@@
$\bullet$ The above argument extends without modification to any countable collection of pairwise distinct points ${x_{i}}_{i\in\mathbb{N}}\subset\mathbb{R}^{d}$. Indeed, $G$ is a countable union of closed subspheres of codimension 1, so by the Baire Category Theorem $\mathbb{S}^{d-1}\setminus G\neq\varnothing$. However, for an uncountable (e.g.,  continuous) data set, one may have $G=\mathbb{S}^{d-1}$, and thus, no separating direction exists.
$\bullet$ If the data set is finite, each $H_{ij}$ is a closed subsphere of dimension $d-2$, and their finite union $G$ is a closed set with empty interior. Hence $G^{c}=\mathbb{S}^{d-1}\setminus G$ is open and dense. In particular, for any $g\in G$ and any $\varepsilon>0$, there exists $g_{\varepsilon}$ in an $\varepsilon-$neighborhood of $g$ such that $g_{\varepsilon}\cdot(x_{i}-x_{j})\neq 0$ for all $i\neq j$, showing that the projection property is stable under small perturbations.
@@
At this point, it becomes evident that at least two hyperplanes are necessary to develop an algorithm for classifying data. Indeed, if we were restricted to using just a single hyperplane, we would be unable to develop a compression process for every data set. Therefore, it is not feasible to classify any $d-$dimensional dataset using a $1$-wide neural network. Thus, $2$ is the minimal width for a neural network (as defined by (1.2)) capable of addressing any classification problem.
@@
Remark 3.2.
The strategy described above has been implemented in Python for a binary classification; the code is available in github.com/Martinshs. This implementation demonstrates how the explicit parameters provided in the proof of Theorem 1.1 can be used to drive each input to its target level set without any training process.
@@ 4.1. Preconditioning of the data:
$w_{1}\cdot x_{i}+b_{1}>0,\quad\forall i\in\llbracket 1,N\rrbracket,$
@@ 4.2.1. The induction strategy
In the sequel we write $\phi(\mathcal{W}^{L},\mathcal{B}^{L},\mathcal{C}_{k})=z_{k}$ when $\phi(\mathcal{W}^{L},\mathcal{B}^{L},x)=z$ for every $x\in\mathcal{C}_{k}$, Therefore, compressing all the classes is equivalent to proving the existence of parameters $\mathcal{W}^{\tilde{L}}$ and $\mathcal{B}^{\tilde{L}}$, $\tilde{L}>0$, and a sequence of different vectors $\{z_{k}\}_{k=0}^{M-1}\subset\mathbb{R}^{2}$ such that
$\phi(\mathcal{W}^{\hat{L}},\mathcal{B}^{\hat{L}},\mathcal{C}_{k})=z_{k},\qquad\text{for every }k\in\llbracket 0,M-1\rrbracket.$ (4.4)
@@
For any $k\in\llbracket 0,M-1\rrbracket$ fixed but arbitrary, we assume that there exist $\tilde{z}_{0}\in\mathbb{R}^{2}$, $\tilde{L}\geq 1$, $\mathcal{W}^{\tilde{L}}$ and $\mathcal{B}^{\tilde{L}}$ such that
$\phi(\mathcal{W}^{\tilde{L}},\mathcal{B}^{\tilde{L}},\mathcal{C}_{k})=\tilde{z}_{0},\qquad\phi(\mathcal{W}^{\tilde{L}},\mathcal{B}^{\tilde{L}},\mathcal{C}\setminus\mathcal{C}_{k})\neq\tilde{z}_{0},$ (4.5)
@@
$\phi(\mathcal{W}^{\tilde{L}},\mathcal{B}^{\tilde{L}},\tilde{z}_{1})\neq\phi(\mathcal{W}^{\tilde{L}},\mathcal{B}^{\tilde{L}},\tilde{z}_{2}),\quad\text{for all }\tilde{z}_{1},\,\tilde{z}_{2}\in\mathcal{C}\setminus\mathcal{C}_{k},\,\tilde{z}_{1}\neq\tilde{z}_{2}.$ (4.6)
@@
$\phi(\mathcal{W}^{L_{1}},\mathcal{B}^{L_{1}},\mathcal{C}_{k})=z_{k},$
@@
In other words, Proposition 4.1 shows that to compress all the classes of points, it is sufficient to compress a single (but arbitrary) class of points without collapsing the points not belonging to that class.
@@
$\bullet$ Data structuring: Construction of hyperplanes driving the data set to some particular structure.
$\bullet$ Compression process: Using the structure established in the prior step, introduce hyperplanes to collapse points belonging to the same class.
@@
$b_{3}^{1}=-\left(\frac{a^{2}_{r_{1}-1}+a^{2}_{r_{1}}}{2}\right),\quad b_{3}^{2}=\frac{a^{2}_{r_{1}}+a^{2}_{r_{1}+1}}{2}$ (4.9)
@@
(2) Inductive Step: The initial step has been achieved. We now aim to show that induction can also be applied successfully. In fact, to do that, it suffices to apply the arguments of the initial step again and again. Note that the model under consideration collapses point to the exact same location, and once this happens, they will never split again in the forthcoming iterations. In this way, if $\alpha_{k}$ denotes the number of elements in $\mathcal{C}_{k}$ i.e. $|\mathcal{C}_{k}|=\alpha_{k}$, all points in $\mathcal{C}_{k}$ will collapse applying $2\alpha_{k}$ times the nonlinear mapping $\sigma$, so that $L_{k}=2\alpha_{k}$.
@@
Data structuring: Denote by $x_{r_{t}}^{2j+1}$ the elements of $\mathcal{C}_{k}^{2j+1}$ for $t\in\llbracket 1,\alpha_{k}\rrbracket$. Observe that, after compressing the first $j$ elements of $\mathcal{C}_{k}^{2j+1}$, we will always have that $x_{r_{t}}^{2j+1}=(0,0)$ for $t\in\llbracket 1,j\rrbracket$ and $x_{r_{j+1}}^{2j+1}=(a_{j+1}^{2j+1},0)$ for some $a_{j+1}^{2j+1}\in\mathbb{R}$. Since we assumed from the beginning that they are not neighboring points of the same class, let $x_{s}^{2j+1}=(a^{2j+1}_{s},0)\notin\mathcal{C}_{k}^{2j+1}$ be the nonzero closest point to $x_{r_{1}}^{2j}=(0,0)$ in the $x-$axis. We also consider $x_{s+1}^{2j+1}=(a_{s+1}^{2j+1},0)$ with $a_{s+1}^{2j+1}\neq 0$, the point to the right of $x_{s}^{2j+1}$ on the $x$-axis (if it does not exist, we take $x_{s+1}^{2j+1}:=x_{s}^{2j+1}+(0.5,0)$).
@@
Compression: We are going to compress $x_{r_{j}}^{2j+2}$ with $x_{r_{j+1}}^{2j+2}$. Let $x_{u}^{2j+2}=(0,a_{u}^{2j+2})$ and $x_{d}^{2j+2}=(0,a_{d}^{2j+2})$ be the points lying above and below $x_{r_{j}}^{2j+2}$ on the $y-$axis, respectively. Also, denote by $x_{rr}^{2j+2}=(a_{rr}^{2j+2},0)$ and $x_{ll}^{2j+2}=(a_{ll}^{2j+2},0)$ the points that are to the right and left of $x_{r_{j+1}}^{2j+2}$ in the $x-$axis. Thus, we consider the parameters